OpenStudy (anonymous):

In a rectangle, the length is twice the width. The diagonal of the rectangle is the square of the width. Find the are of the rectangle.

OpenStudy (phi):

 Area= L*W L= 2W (L is twice the width) so, Area= 2W^2 Also, we know the diagonal D= W^2. But by the pythagorean theorem, D^2= L^2 + W^2 so L^2 + W^2=W^4 substitute for L=2W to get 4W^2+W^2= W^4 divide by W^2 to get 4+1= W^2 or W^2 =5 remember that Area= 2W^2 (see above) or Area= 2*5=10